Assume negligible ionization to determine the percent ionization of each of the following acids and then indicate whether the assumption was valid.

a) 0.800 M solution of cyanic acid (HCNO, pKa = 3.46)

% ionization =   Valid: Yes or No

b) 0.1000 M solution of hydroazoic acid (HN3, pKa = 4.60)

% ionization =   Valid: Yes or No

c) 0.300 M solution of benzoic acid (HC7H5O2, pKa = 4.19)

% ionization =   Valid: Yes or No
Answer #1

a) % ionization = 2.08 %

Valid : Yes

b) % ionization = 1.58 %

Valid : Yes

c) % ionization = 1.47 %

Valid : Yes

Explanation

(a) pKa = 3.46

Ka = 10-pKa

Ka = 10-3.46

Ka = 3.467 x 10-4

[H+] = (Ka * C)1/2

[H+] = (3.467 x 10-4 * 0.800 M)1/2

[H+] = 0.016655 M

percent ionization = ([H+] / C) * 100

percent ionization = (0.016655 M / 0.800 M) * 100

percent ionization = 2.08 %

Since percent ionization is less than 5%, therefore, assumption is valid.